A cash register contains only 10 dollars and fifty dollars. It contains five times as many as many 10 dollar bills as fifty dollar bills and the total amount is 1800. How many fifty dollar bills are in the cash register?

Answer by Boreal(15235)

$10=x
$50=(1/5)x
10x + 50(x/5)=1800
20x=1800; 50 * (1/5)x=10x
x=90
There are 90 $10 bills for $900
There are 18 $50 bills for $900